# Two-interface and thin filament approximation in Hele–Shaw channel flow

### v2

@@ 1 Introduction
The two-interface Hele–Shaw model in the limit that the interfaces are close is of great importance as even if the fluid region is not initially thin, the effect of the Saffman–Taylor instability will result in a thin fluid region or filament developing (see for example the experimental results in Ward & White (2011); Morrow et al. (2023), and Cardoso & Woods (1995), although the latter study focuses on a much less viscous thin annular region in a more viscous ambient fluid). This formation of a thin filament precedes the fluid ‘bursting’, at which point the two inviscid regions meet and the pressure rapidly equalises. The breakup of a thin viscous filament in a Hele–Shaw cell (with surface tension but in the absence of a driving pressure difference) has also been examined by the use of the lubrication approximation (Almgren, 1996; Almgren et al., 1996; Constantin et al., 1993; Dupont et al., 1993; Goldstein et al., 1993, 1995, 1998). The complicated, but self-similar break-up behaviour of the filament in particular (where the filament thickness goes to zero at a finite time and point in space) is detailed in Almgren et al. (1996).
@@
In this article we consider two-interface Hele–Shaw flow in a channel, including the effects of both driving pressure difference and surface tension, with particular focus on the case in which the fluid interfaces become close together. In Section 2 we describe the full two-interface model and its stability. In Section 3 we derive a simplified model (the thin filament approximation) that applies when the two interfaces are close together, by applying the lubrication approximation (up to second order) in a coordinate system following the filament centreline. This approximation represents a generalisation of the model by Farmer & Howison (2006) by including both the effects of surface tension, as well as higher order terms, which include the effect of flow along the filament. In section 4 we demonstrate the importance of including these higher order terms by showing (both numerically and by constructing similarity solutions) that the leading order problem generically blows up in finite time, even in the presence of surface tension. We also find quasi-travelling wave solutions, which may be thought of as the analogue of Saffman–Taylor fingers, which generalise the ‘grim reaper’ exact solution found by Farmer & Howison (2006). In Section 5 we compute numerical solutions of the thin filament model including the higher order regularising terms. Solutions of the original two-interface model, found using a level set method, are seen to approach the thin filament model results as the resolution of the level set method is increased. We show the general behaviour of our thin filament model is not to tend toward a quasi-travelling wave solution, but instead develop a rapidly expanding ‘bubble’ of circular shape and decreasing thickness.
@@ 2.1 Hele–Shaw flow equations
We consider a Hele–Shaw channel of nondimensional width $2\pi$ in the $x$-direction. The fluid region $\Omega$ is bounded above and below by interfaces $\partial\Omega_{\mathrm{U}}$ and $\partial\Omega_{\mathrm{L}}$, respectively. A nondimensional pressure difference $P$ acts to push the fluid region in the positive $y$-direction, while surface tension acts on both interfaces (see Fig. 1). The standard governing equations in nondimensional form are (Morrow et al., 2021, e.g.)
@@
with $\phi$ the velocity potential, $v_{\mathrm{n}}$ the normal velocity of each interface, $\kappa_{\mathrm{U}}$ and $\kappa_{\mathrm{L}}$ the curvatures of each interface, $P$ the nondimensional pressure difference, and $\sigma$ the nondimensional surface tension. The normal vector $\hat{\bm{n}}$ of each interface is defined so that a flat interface has normal in the positive $y$-direction, and the signs of curvatures $\kappa$ are chosen such that positive $\kappa$ implies a concave interface in the positive $y$-direction (for example, both interfaces in Fig. 1 have negative curvature at $x=0$ and positive curvature at $x=\pm\pi$). This choice of sign convention for both interfaces will simplify the thin filament derivation in the next section. While one of the parameters $P$ or $\sigma$ (if nonzero) could be scaled out by choosing an appropriate time scale, retaining both parameters aids in understanding the effects of the terms that arise in our later analysis.
For definiteness, we consider flow in a channel of nondimensional width $2\pi$, with no-flux boundary conditions on the channel walls:
@@ 2.2 Linear stability
The two-interface Hele–Shaw configuration has an exact base state where both interfaces are horizontal, and the fluid region moves upward at constant velocity $v_{\mathrm{n}}=P/h_{0}$, where $h_{0}$ is the distance between the two interfaces. To examine the stability of this configuration, write the system (2.1) in Cartesian ($x,y$) coordinates, and define the upper and lower interfaces as $y=f_{\mathrm{U}}(x,t)$ and $y=f_{\mathrm{L}}(x,t)$, respectively. In addition to Laplace’s equation for $\phi$, the boundary conditions (1b)-(1d) are
@@ 3 Thin filament approximation
In this section we derive an approximation of the two-interface flow by considering the thickness of the fluid region (that is, the distance between the two interfaces) to be small. We will start by converting the governing equations and boundary conditions into an intrinsic coordinate system $(x,y)\mapsto(s,n)$ in which the filament is represented by a centreline curve $\bar{\bm{x}}(s,t)=(\bar{x},\bar{y})$, with $s$ the arclength coordinate, and $n$ the coordinate normal to this centreline; thus
@@
(see Fig. 3). Here $\bm{n}=(-\bar{y}_{s},\bar{x}_{s})$ is the centreline normal (again we use the variable subscript $s$ to represent the $s$ partial derivative, for brevity). In this coordinate system we specify the upper and lower interfaces to occur at $n=\pm h(s,t)/2$, respectively, so that the normal thickness is $h$. The upper interface is thus given by the curve
@@
$\frac{D\bm{x}_{\mathrm{U}}}{Dt}=\frac{D\bm{x}}{Dt}+\frac{h}{2}\frac{D\bm{n}}{Dt}+\frac{1}{2}\frac{Dh}{Dt}\bm{n}.$
@@
The expressions (11) and (13) are exact, but we have not yet determined the potential $\phi$. We now formally take the lubrication approximation by substituting $n=\epsilon N$, $h=\epsilon H$, and $t=\epsilon T$, taking $\epsilon\ll 1$. Unlike most standard lubrication models where the leading order behaviour in $\epsilon$ is all that is required, in our case we will see that terms at order $\epsilon^{2}$ are necessary to fully regularise the instability, and so we will expand to this order.
@@
$\frac{DH}{DT}=\kappa Hv_{\mathrm{n}}-\epsilon^{2}\frac{\partial}{\partial s}\left[-\frac{H^{3}\kappa_{s}V_{0}}{12}-\frac{H^{2}H_{s}\kappa V_{0}}{4}+H\sigma\kappa_{1s}\right]$ (15c)
In (15b), (15c), the order of terms in the lubrication parameter $\epsilon$ is explicit. We can of course remove the scaling by returning to $h,t$ variables, resulting in
@@
with $v_{0}=(P+2\sigma\kappa)/h$ the leading-order velocity, and $\kappa_{1,2}$ equal to $K_{1,2}$ in (14) with $H\mapsto h$.
As will be further detailed in section 4, the higher order terms have been included in the above derivation as they are needed to fully regularise the instability. In particular, the higher spatial derivative of thickness $h$ resulting from the interface curvature correction term $\kappa_{1}$ plays a crucial role. In the case where driving pressure $P$ and centreline curvature $\kappa$ both vanish, (3) reduces to the well-known thin film equation
@@
where $h=h(s,t)$, and arclength $s$ becomes time-independent. This equation has been studied extensively in the context of droplet breakup in Hele–Shaw flow (Almgren, 1996; Almgren et al., 1996; Constantin et al., 1993; Dupont et al., 1993; Goldstein et al., 1993, 1995, 1998).
@@ 3.1 Stability
As with the full two-interface model (2.1), the thin filament approximation (3) has an exact solution comprising a straight filament of uniform thickness $h_{0}$ moving upward with speed $P/h_{0}$. To test the stability of this straight filament, we write the centreline $\bar{\bm{x}}(s,t)$ in Cartesian coordinates as $y=f(x,t)$, write thickness $h$ as a function of $x$ and $t$, and perturb the straight filament:
$h(x,t)=h_{0}+\tilde{h}\mathrm{e}^{\mathrm{i}kx+\lambda t},\qquad f(y,t)=\frac{P}{h_{0}}t+\tilde{f}\mathrm{e}^{\mathrm{i}kx+\lambda t}.$
@@
We note that the linear stability of the filament model (3) does not agree exactly with the full model, even for zero surface tension. When $\sigma=0$, the error in the filament model is due to the final term $P^{2}h_{0}^{2}k^{6}/144$, which arises from the product of the two higher order correction terms in the off-diagonal terms in the determinant. This term is (implicitly) of order $\epsilon^{4}$, which is of higher order than the model is accurate to.
@@ 4.1 The unregularised (zero surface tension) leading-order model
The previous two exact solutions show only a subset of the possible singular behaviours of the ill-posed zero-surface tension model; as centrelines are given by level curves of a harmonic function $t=\Im(w(z))$, with thickness given by (22), it is likely that solutions may be constructed that exhibit thickness becoming infinite, self-intersection of the interface, and so on. As the ill-posed ($\sigma=0$) problem is not characteristic of the full problem, we do not go into further details here.
Figure 4: Exact solutions to the thin filament equation in the absence of surface tension. (a) the solution (24) that evolves from an initial condition that results in corner formation with the generic angle of $\pi/2$, and (b) the solution (25) that evolves from an initial condition that results in corner formation with a non-generic angle of $\pi/3$. The parameter values are $a=0.1$, $\delta=0.2$. In both examples, the centrelines are plotted in solid blue while the black dashed lines are the upper and lower interfaces (found by adding or subtracting half the thickness $h$ in the normal direction).
@@ 4.2 Singularity formation in the leading-order model with surface tension
where $C$ is an arbitrary scaling factor. Expanding near the singular point results in $\gamma=\frac{n}{n-1}>1$, which is consistent with the assumptions made on the similarity exponents above.
The equation for $u$ can be solved implicitly by letting $u$ be the independent variable, which allows us to construct parametric solutions for $\hat{F}$ and $\hat{G}$:
@@ 4.3 Quasi-travelling wave solutions
where $t_{0}$ is a finite time. A solution of this form is similar to, but not exactly, a travelling wave, as the centreline has a fixed shape but moves to infinity (with speed unbounded) as $t\to t_{0}$, while the thickness linearly decreases to zero. The parameter $B$ is the analogue of a travelling wave speed. Solutions to (26) of this form generalise the ‘grim-reaper’ solutions to the zero surface tension problem found in Farmer & Howison (2006), and are also candidates for asymptotically valid solutions to the higher-order system (3), since as the thickness goes to zero, we would expect the higher-order terms vanish at faster rate than the leading-order terms. We will see in Section 5 that these quasi-travelling waves do not appear to be attractors, but will compute them here for completeness.
@@
$-\left[\hat{F}^{\prime}\left(1+\epsilon\frac{\hat{F}^{\prime\prime}}{(1+\hat{F}^{\prime 2})^{3/2}}\right)\right]^{\prime}=(1+\hat{F}^{\prime 2})\left(1+\hat{\sigma}\frac{\hat{F}^{\prime\prime}}{(1+\hat{F}^{\prime 2})^{3/2}}\right),\qquad\hat{\sigma}=\frac{2\sigma}{BP}.$ (30)
@@
For a semi-infinite curve, the appropriate interval is $-\pi/2<\theta<\pi/2$, where the nose at $\theta=0$ is a singular point, at which we require $K(0)=-1$.
@@
$\frac{\mathrm{d}\hat{x}}{\mathrm{d}\theta}=-\frac{\cos\theta}{K},\qquad\frac{\mathrm{d}\hat{F}}{\mathrm{d}\theta}=\frac{\sin\theta}{K}.$
@@
$K=K_{\infty}=-\cot\theta\log\left(\frac{\cos(\theta/2)+\sin(\theta/2)}{\cos(\theta/2)-\sin(\theta/2)}\right).$
Plots of these quasi-travelling wave solutions are shown in Fig. 6b. The effect of $\hat{\sigma}$, and thus of surface tension, is weak, as all solutions are bounded between the $\hat{\sigma}=0$ and $\hat{\sigma}\to\infty$ limits. The scaled curvature at the nose is required to be $-1$ in all cases, and thus returning to the unscaled system, we expect the curvature at the nose to be $-1/B$, that is, inversely proportional to the wave speed parameter. We will see in Section 5, however, that these quasi-travelling wave solutions do not act as attractors for a generic initial value problem.

### v3

@@ 1 Introduction
The two-interface Hele–Shaw model in the limit that the interfaces are close is of great importance as, even if the fluid region is not initially thin, the effect of the Saffman–Taylor instability will result in a thin fluid region or filament developing (see for example the experimental results in Ward & White (2011); Morrow et al. (2023), and Cardoso & Woods (1995), although the latter study focuses on a much less viscous thin annular region in a more viscous ambient fluid). This formation of a thin filament precedes the fluid ‘bursting’, at which point the two inviscid regions meet and the pressure rapidly equalises. The breakup of a thin viscous filament in a Hele–Shaw cell (with surface tension but in the absence of a driving pressure difference) has also been examined by the use of the lubrication approximation (Almgren, 1996; Almgren et al., 1996; Constantin et al., 1993; Dupont et al., 1993; Goldstein et al., 1993, 1995, 1998). The complicated, but self-similar break-up behaviour of the filament in particular (where the filament thickness goes to zero at a finite time and point in space) is detailed in Almgren et al. (1996).
@@
In this article we consider two-interface Hele–Shaw flow in a channel, including the effects of both driving pressure difference and surface tension, with particular focus on the case in which the fluid interfaces become close together. In Section 2 we describe the full two-interface model and its stability. In Section 3 we derive a simplified model (the thin filament approximation) that applies when the two interfaces are close together, by applying the lubrication approximation (up to second order) in a coordinate system following the filament centreline. The application of lubrication theory in an evolving, curvlinear coordinate system is similar to that applied in Van De Fliert et al. (1995) and Howell (2003), although taking such an approximation to higher order is not standard. Our approximation also represents a generalisation of the model by Farmer & Howison (2006) by including both the effects of surface tension, as well as higher order terms, which include the effect of flow along the filament. In section 4 we demonstrate the importance of including these higher order terms by showing (both numerically and by constructing similarity solutions) that the leading order problem generically blows up in finite time, even in the presence of surface tension. We also find quasi-travelling wave solutions, which may be thought of as the analogue of Saffman–Taylor fingers, which generalise the ‘grim reaper’ exact solution found by Farmer & Howison (2006). In Section 5 we compute numerical solutions of the thin filament model including the higher order regularising terms. Solutions of the original two-interface model, found using a level set method, are seen to approach the thin filament model results as the resolution of the level set method is increased. We show the general behaviour of our thin filament model is not to tend toward a quasi-travelling wave solution, but instead develop a rapidly expanding ‘bubble’ of circular shape and decreasing thickness.
@@ 2.1 Hele–Shaw flow equations
We consider flow in a Hele–Shaw channel of nondimensional width $2\pi$ in the $x$-direction. The fluid region $\Omega$ is bounded above and below by interfaces $\partial\Omega_{\mathrm{U}}$ and $\partial\Omega_{\mathrm{L}}$, respectively. A nondimensional pressure difference $P$ acts to push the fluid region in the positive $y$-direction, while surface tension acts on both interfaces (see Fig. 1). The standard governing equations in nondimensional form are (Morrow et al., 2021, e.g.)
@@
with $\phi$ the velocity potential, $v_{\mathrm{n}}$ the normal velocity of each interface, $\kappa_{\mathrm{U}}$ and $\kappa_{\mathrm{L}}$ the curvatures of each interface, $P$ is the nondimensional pressure difference, and $\sigma$ the nondimensional surface tension. In (1) the normal vector $\hat{\bm{n}}$ of each interface is defined so that a flat interface has normal in the positive $y$-direction, and the signs of curvatures $\kappa$ are chosen such that positive $\kappa$ implies a concave interface in the positive $y$-direction (for example, both interfaces in Fig. 1 have negative curvature at $x=0$ and positive curvature at $x=\pm\pi$). This choice of sign convention for both interfaces will simplify the thin filament derivation in the next section.
For definiteness, we consider no-flux boundary conditions on the channel walls:
@@
The system (1) is derived from the dimensional system by introducing the length scale $[x]$, the pressure scale $[p]$, and defining the dimensionless velocity potential $\phi$ in terms of the pressure $p$:
$[x]=\frac{\ell}{2\pi},\qquad[p]=\frac{12\mu[x]^{2}}{b^{2}[t]},\qquad\phi=\frac{p_{U}-p}{[p]},$
so that
$P=\frac{p_{U}-p_{L}}{[p]}\qquad\sigma=\frac{\hat{\gamma}}{[p][x]},$
where $p_{U}$ and $p_{L}$ are the upper and lower inviscid fluid pressures, $b$ is the plate separation, $\ell$ is the channel width, $\mu$ the viscosity, $\hat{\gamma}$ is the surface tension, and $[t]$ is an arbitrary time scale. While one of the parameters $P$ or $\sigma$ (if nonzero) could be scaled out by choosing an appropriate time scale, retaining both parameters aids in understanding the effects of the terms that arise in our later analysis.
@@ 2.2 Linear stability
The two-interface Hele–Shaw configuration has an exact base state where both interfaces are horizontal, and the fluid region moves upward at constant velocity $v_{\mathrm{n}}=P/h_{0}$, where $h_{0}$ is the distance between the two interfaces. To examine the stability of this configuration, write the system (1) in Cartesian ($x,y$) coordinates, and define the upper and lower interfaces as $y=f_{\mathrm{U}}(x,t)$ and $y=f_{\mathrm{L}}(x,t)$, respectively. In addition to Laplace’s equation for $\phi$, the boundary conditions (1b)-(1d) are
@@ 3 Thin filament approximation
3.1 Model in intrinsic coordinate system
In this section we derive an approximation of the two-interface flow by considering the thickness of the fluid region (that is, the distance between the two interfaces) to be small. This approximation will be valid when the separation of the two interfaces is smaller than the radius of curvature of either interface, but still larger than the plate separation (if the interface separation is the same order as the plate separation, the depth averaging that leads to two-dimensional Hele–Shaw flow (1) is no longer valid).
We will start by converting the governing equations and boundary conditions into an intrinsic coordinate system $(x,y)\mapsto(s,n)$ in which the filament is represented by a centreline curve $\bar{\bm{x}}(s,t)=(\bar{x},\bar{y})$, with $s$ the arclength coordinate, and $n$ the coordinate normal to this centreline; thus
@@ 3.1 Model in intrinsic coordinate system
(e.g., Van De Fliert et al., 1995; Howell, 2003). The coordinate system is represented schematically in Fig. 3. Here $\bm{n}=(-\bar{y}_{s},\bar{x}_{s})$ is the centreline normal (again we use the variable subscript $s$ to represent the $s$ partial derivative, for brevity). In this coordinate system we specify the upper and lower interfaces to occur at $n=\pm h(s,t)/2$, respectively, so that the normal thickness is $h$. The upper interface is thus given by the curve
@@
$\frac{D\bm{x}_{\mathrm{U}}}{Dt}=\frac{D\bar{\bm{x}}}{Dt}+\frac{h}{2}\frac{D\bm{n}}{Dt}+\frac{1}{2}\frac{Dh}{Dt}\bm{n}.$
@@
The expressions (11) and (13) are exact, but we have not yet determined the potential $\phi$. We now formally take the lubrication approximation by substituting $n=\epsilon N$, $h=\epsilon H$, and $t=\epsilon T$, taking $\epsilon\ll 1$. In this expansion we assume the curvature $\kappa=O(1)$ (or smaller), and the surface tension $\sigma=O(1)$. Unlike most standard lubrication models where the leading order behaviour in $\epsilon$ is all that is required, in our case we will see that terms at order $\epsilon^{2}$ are necessary to fully regularise the instability, and so we will expand to this order.
@@
$\frac{DH}{DT}=\kappa HV_{\mathrm{n}}-\epsilon^{2}\frac{\partial}{\partial s}\left[-\frac{H^{3}\kappa_{s}V_{0}}{12}-\frac{H^{2}H_{s}\kappa V_{0}}{4}+H\sigma K_{1s}\right]$ (15c)
In (15b), (15c), the order of terms in the lubrication parameter $\epsilon$ is explicit. We can of course formally remove the scaling by returning to $h,t$ variables, resulting in
@@
with $v_{0}=(P+2\sigma\kappa)/h$ the leading-order velocity, and $\kappa_{1}=\epsilon^{-1}K_{1}$, $\kappa_{2}=\epsilon^{-2}K_{2}$ the rescaled versions of (14):
$\kappa_{1}=\frac{h\kappa^{2}}{2}+\frac{h_{ss}}{2},\qquad\kappa_{2}=\frac{h^{2}\kappa^{3}}{4}+\frac{hh_{s}\kappa_{s}}{4}+\frac{hh_{ss}\kappa}{2}+\frac{h_{s}^{2}\kappa}{8}.$ (16c)
The higher order terms have been included in the above derivation as they are needed to fully regularise the instability. This will become apparent when we examine the behaviour of the leading order system in Section 4. In particular, the higher spatial derivative of thickness $h$ resulting from the interface curvature correction term $\kappa_{1}$ plays a crucial role. In the case where driving pressure $P$ and centreline curvature $\kappa$ both vanish, (16) reduces to the well-known thin film equation
@@
where $h=h(s,t)$, and arclength $s$ becomes time-independent. This equation has been studied extensively in the context of droplet breakup in Hele–Shaw flow (Almgren, 1996; Almgren et al., 1996; Constantin et al., 1993; Dupont et al., 1993; Goldstein et al., 1993, 1995, 1998). In our case, this fourth order spatial derivative term stabilises high wavenumber perturbations, as is shown in the following stability analysis of the filament model.
@@ 3.2 Stability
As with the full two-interface model (1), the thin filament approximation (16) has an exact solution comprising a straight filament of uniform thickness $h_{0}$ moving upward with speed $P/h_{0}$. To test the stability of this straight filament, we write the centreline $\bar{\bm{x}}(s,t)$ in Cartesian coordinates as $y=f(x,t)$, write thickness $h$ as a function of $x$ and $t$, and perturb the straight filament:
$h(x,t)=h_{0}+\tilde{h}\mathrm{e}^{\mathrm{i}kx+\lambda t},\qquad f(x,t)=\frac{P}{h_{0}}t+\tilde{f}\mathrm{e}^{\mathrm{i}kx+\lambda t}.$
@@
We note that the linear stability of the filament model (16) does not agree exactly with the full model, even for zero surface tension. When $\sigma=0$, the error in the filament model is due to the final term $P^{2}h_{0}^{2}k^{6}/144$, which arises from the product of the two higher order correction terms in the off-diagonal terms in the determinant. This term is (implicitly) of order $\epsilon^{4}$, which is of higher order than the model is accurate to. If further terms were taken in the lubrication expansion, they would contribute further order $\epsilon^{4}$ terms which would presumably cancel with the term present here.
@@ 4.1 The unregularised (zero surface tension) leading-order model
The previous two exact solutions show only a subset of the possible singular behaviours of the ill-posed zero-surface tension model; as centrelines are given by level curves of a harmonic function $t=\Im(w(z))$, with thickness given by (22), a variety of other solution behaviour is possible. One such possibility is a zero-angle cusp that results from a square root singularity, an example of which is the complex function
$A+\mathrm{i}t=w(z)=\delta\left(z-a\mathrm{i}\sqrt{1-\mathrm{e}^{-\mathrm{i}z}}\right).$ (26)
This function has a square root singularity at $z=0$, as well as the appropriate periodicity. The level curve that passes through this singularity occurs at $t=0$, so as $t\to 0^{-}$, a zero angle, backward-facing cusp forms on the interface. The thickness $h$, which is given by $|w^{\prime}(z)|$, becomes infinite at the cusp, so that the interface at the cusp becomes stationary. The profiles and thickness corresponding to (26) are shown in Fig. 4c–d, respectively, for the parameter values of $\delta=0.2$ and $a=0.5$.
Figure 4: Exact solutions to the thin filament equation in the absence of surface tension. (a) the solution (24) that evolves from an initial condition that results in corner formation with the generic angle of $\pi/2$, and (b) the solution (25) that evolves from an initial condition that results in corner formation with a non-generic angle of $\pi/3$. The parameter values are $a=0.1$, $\delta=0.2$. In both of these examples, the centrelines are plotted in solid blue while the black dashed lines are the upper and lower interfaces (found by adding or subtracting half the thickness $h$ in the normal direction). (c), the profiles, and (d) the thickness, for the example (26) with $\delta=0.2$ and $a=0.5$, which forms a zero angle cusp with infinite thickness.
@@ 4.2 Singularity formation in the leading-order model with surface tension
where $C$ is an arbitrary scaling factor. Expanding near the singular point results in $\gamma=n/(n-1)>1$, which is consistent with the assumptions made on the similarity exponents above.
Given $\gamma=n/(n-1)$, the equation for $u$ can be solved implicitly by letting $u$ be the independent variable, which allows us to construct parametric solutions for $\hat{F}$ and $\hat{G}$:
@@ 4.3 Quasi-travelling wave solutions
where $t_{0}$ is a finite time. A solution of this form is similar to, but not exactly, a travelling wave, as the centreline has a fixed shape but moves to infinity (with speed unbounded) as $t\to t_{0}$, while the thickness linearly decreases to zero. The parameter $B$ is the analogue of a travelling wave speed. Solutions to (27) of this form generalise the ‘grim-reaper’ solutions to the zero surface tension problem found in Farmer & Howison (2006), and are also candidates for asymptotically valid solutions to the higher-order system (16), since as the thickness goes to zero, we would expect the higher-order terms to vanish at faster rate than the leading-order terms. We will see in Section 5 that these quasi-travelling waves do not appear to be attractors, but will compute them here for completeness.
@@
$-\left[\hat{F}^{\prime}\left(1+\hat{\sigma}\frac{\hat{F}^{\prime\prime}}{(1+\hat{F}^{\prime 2})^{3/2}}\right)\right]^{\prime}=(1+\hat{F}^{\prime 2})\left(1+\hat{\sigma}\frac{\hat{F}^{\prime\prime}}{(1+\hat{F}^{\prime 2})^{3/2}}\right),\qquad\hat{\sigma}=\frac{2\sigma}{BP}.$ (31)
@@
For a semi-infinite curve, the appropriate interval is $-\pi/2<\theta<\pi/2$, where the nose at $\theta=0$ is a singular point, at which we require $K(0)=-1$. In the limit of the tail of the travelling wave $\theta\to\pi/2$, we must have $K\sim\theta-\pi/2\to 0$.
@@
$\frac{\mathrm{d}\hat{x}}{\mathrm{d}\theta}=-\frac{\cos\theta}{K},\qquad\frac{\mathrm{d}\hat{F}}{\mathrm{d}\theta}=\frac{\sin\theta}{K}.$ (33)
@@
$K=K_{\infty}=-\cot\theta\log\left(\frac{\cos(\theta/2)+\sin(\theta/2)}{\cos(\theta/2)-\sin(\theta/2)}\right).$ (34)
Plots of these quasi-travelling wave solutions are shown in Fig. 6b. The effect of $\hat{\sigma}$, and thus of surface tension, is weak, as all solutions are bounded between the $\hat{\sigma}=0$ and $\hat{\sigma}\to\infty$ limits. The scaled curvature at the nose is required to be $-1$ in all cases, and thus returning to the unscaled system, we expect the curvature at the nose to be $-1/B$, that is, inversely proportional to the wave speed parameter.
The corresponding scaled thickness in the $y$ direction is $\bar{H}=\sec^{2}\theta(1+\hat{\sigma}\kappa)$. The time-dependent thickness in the centreline-normal direction is thus
$h=\frac{P(t_{0}-t)}{B\cos\theta}(1+\hat{\sigma}\kappa).$
At a fixed time $t_{0}$ then, the thickness is unbounded in the tails as $\theta\to\pm\pi/2$. If, however, we consider a fixed value of $y$ (in the non-travelling coordinate system) as the singular time $t_{0}$ is approached, then from (30) and (33), the corresponding value of $\theta$ behaves as
$\pi/2-\theta\sim(t_{0}-t)\mathrm{e}^{y/B},\qquad t\to t_{0}.$
Thus $h=O(\mathrm{e}^{-y/B})$ is bounded for a fixed value of $y$ as $t\to t_{0}$. The exponential decay in $y$ of the thickness also means that a quasi-travelling wave solution does not require infinite mass to form, even as it length is unbounded in the $y$-direction as $t\to t_{0}$.
In Section 5 we will simulate the second order thin filament model, including with initial conditions close to a quasi-travelling wave solution. These simulations will demonstrate that these quasi-travelling wave solutions do not appear to be stable.